OpenStudy (anonymous):

Find the coordinates of the midpoint of the segment connecting P(6, –6) and Q(–2, 7). (4, 1)......??

OpenStudy (hoblos):

x= (6-2)/2 = 2 y= ( -6+7)/2 = 1/2 (2,1/2)

OpenStudy (anonymous):

\[(\frac{6-2}{2},\frac{-6+7}{2})\]

OpenStudy (anonymous):

2,1/2